1. Find the area of a sector of a circle with radius 6 cm if angle of the sector is 60°.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
It is given that the angle of the sector is 60° We know that the area of sector = (θ/360°)×πr2 ∴ Area of the sector with angle 60° = (60°/360°)×πr2 cm2 = (36/6)π cm2 = 6×22/7 cm2 = 132/7 cm2